# SBP-YOLO:A Lightweight Real-Time Model for Detecting Speed Bumps and Potholes

### v2

@@ Abstract
Reliable and real-time detection of road speed bumps and potholes is crucial for anticipatory perception in advanced suspension systems, enabling timely and adaptive damping control. However, achieving high accuracy and efficiency on embedded platforms remains challenging due to limited computational resources and the small scale of distant targets. This paper presents SBP-YOLO, a lightweight and high-speed detection framework tailored for bump and pothole recognition. Based on YOLOv11n, the model integrates GhostConv and VoVGSCSPC modules into the backbone and neck to reduce computation while enhancing multi-scale semantic features. To improve small-object detection, a P2-level branch is introduced, with a Lightweight and efficient detection head (LEDH) mitigating the added computational overhead without compromising accuracy. A hybrid training strategy combining NWD loss, BCKD-based knowledge distillation, and Albumentations-driven augmentation further enhances localization precision and robustness. Experiments show that SBP-YOLO achieves 87.0% mAP, outperforming the YOLOv11n baseline by 5.8%. After TensorRT FP16 quantization, it runs at 139.5 FPS on Jetson AGX Xavier, delivering a 12.4% speedup over the P2-enhanced YOLOv11. These results validate the effectiveness of the proposed method for fast and low-latency road condition perception in embedded suspension control systems.
@@ 1 Introduction
Adjustable suspension systems, such as magnetorheological semi-active [1, 2] and fully active suspensions [3], offer real-time tunability of damping and stiffness to enhance ride comfort and handling stability [4]. However, their performance largely depends on timely and accurate perception of upcoming road conditions [5]. In the absence of reliable preview information, these systems are limited to reactive control, making it difficult to fully exploit their potential in suppressing body vibrations and improving vehicle dynamics. Among various road disturbances, speed bumps and potholes represent common yet critical anomalies that directly impact ride quality [6]. Early detection of such features enables predictive suspension control, allowing damping strategies to be adjusted in advance[7, 8, 9]. Existing methods for road anomaly detection rely on vehicle dynamics, vibration analysis, or multi-sensor fusion (e.g., accelerometers, gyroscopes, and GPS) [10, 11, 12]. While effective, these approaches are often task-specific, computationally intensive, or limited by high hardware costs, making them unsuitable for real-time embedded deployment in advanced suspension systems.
@@
Camera-based visual perception offers a cost-effective and information-rich alternative to traditional sensor-based approaches for road condition monitoring. Deep learning has substantially advanced vision-based anomaly detection by enabling proactive identification of road defects such as speed bumps and potholes. Compared to acceleration signal analysis, vision-based methods provide earlier and more reliable detection, offering critical input for real-time suspension control. The YOLO (You Only Look Once) series has further improved the trade-off between detection accuracy and inference speed [13]. Bučko et al. [14] evaluated YOLOv3 for pothole detection under diverse lighting and weather conditions, and compared it with Sparse R-CNN. Their results showed that while Sparse R-CNN offered slightly higher robustness in extreme visual degradation, YOLOv3 remained competitive due to its lower hardware requirements and stable performance across most conditions. Subsequent versions introduced structural improvements: YOLOv4 adopted CSPDarknet and PANet for enhanced feature aggregation, while its lightweight variant YOLOv4-tiny was deployed by Park et al. [15] for pothole detection with superior real-time capability. ASAD et al. [16] implemented a lightweight YOLOv4-tiny model on a Raspberry Pi device for edge-based pothole detection, achieving a detection accuracy of 90% at 31.76 FPS. YOLOv5 [17] incorporated anchor-free heads and auto-learned anchor boxes, whereas YOLOv8 [18] leveraged decoupled heads and improved data augmentation techniques. Khan et al. [19] further optimized YOLOv8 with customized augmentations for autonomous driving scenarios. Building on these, Bhavana et al. [20] proposed POT-YOLO based on YOLOv8, achieving 97.6% accuracy in detecting complex pavement defects. More recently, YOLOv11 [21] introduced the C3k2 residual block and C2PSA attention mechanism to enhance multiscale feature representation. Its modular architecture spans from lightweight to large-scale variants, supporting flexible deployment across both embedded and high-performance platforms.
Despite progress in vision-based detection of potholes and speed bumps, real-world use under dynamic driving remains difficult. At long range, these anomalies appear small, making them hard to detect. Motion blur from fast driving degrades features, while shadows, glare, and dim lighting obscure road cues and reduce robustness. Limited computing power on embedded systems also limits input size and model depth, making real-time use harder. To address these challenges, this paper proposes a lightweight and efficient model tailored for speed bump and pothole detection. The main contributions are as follows:
  • $\bullet$
@@
    A lightweight and efficiency detection head (LEDH) is introduced to reduce the computational overhead of the P2 layer, maintaining accuracy for small and distant targets. Combined with GhostConv modules and VoVGSCSPC blocks in the backbone and neck, the proposed architecture enhances feature extraction efficiency and robustness under diverse road and environmental conditions.
  • $\bullet$
@@
    A comprehensive training strategy is adopted to improve detection performance in dynamic scenarios, incorporating NWD Loss [22] for precise localization, BCKD [23] for knowledge distillation, and Albumentations-based [24] data augmentation to simulate motion blur, lighting variations, and adverse weather conditions.
  • $\bullet$
@@
    The model is quantized to FP16 and optimized using TensorRT, achieving an inference speed of 139.5 FPS on the NVIDIA Jetson AGX Xavier, thereby meeting real-time requirements for embedded road condition detection.
@@ 2.2 GhostConv Module
Under typical settings (e.g., $c_{1}\in[64,256]$, $k_{m}=3$, $k_{c}=5$), GhostConv reduces parameters and FLOPs to approximately 48%–55% of those of standard convolution. This demonstrates its effectiveness for efficient deployment in embedded detection frameworks. Moreover, Fig. 2(c) shows visualized feature maps from the 5th and 7th layers in the backbone. The GhostConv-enhanced model exhibits stronger responses to distant small-scale targets, highlighting its improved spatial awareness in complex road scenes.

### v4

@@ Abstract
Speed bumps and potholes are the most common road anomalies, significantly affecting ride comfort and vehicle stability. Preview-based suspension control mitigates their impact by detecting such irregularities in advance and adjusting suspension parameters proactively. Accurate and real-time detection is essential, but embedded deployment is constrained by limited computational resources and the small size of targets in input images. To address these challenges, this paper proposes SBP-YOLO, an efficient detection framework for speed bumps and potholes in embedded systems. Built upon YOLOv11n, it integrates GhostConv and VoVGSCSPC modules in the backbone and neck to reduce computation while enhancing multi-scale semantic features. A P2-level branch improves small-object detection, and a lightweight and efficient detection head (LEDH) maintains accuracy with minimal overhead. A hybrid training strategy further enhances robustness under varying road and environmental conditions, combining NWD loss, BCKD knowledge distillation, and Albumentations-based augmentation. Experiments show that SBP-YOLO achieves 87.0% mAP, outperforming the YOLOv11n baseline by 5.8%. After TensorRT FP16 quantization, it runs at 139.5 FPS on Jetson AGX Xavier, yielding a 12.4% speedup over the P2-enhanced YOLOv11. These results demonstrate the framework’s suitability for fast, low-latency road condition perception in embedded suspension control systems.
Keywords Road Anomaly Detection, Lightweight Real-Time Object Detection, Embedded Vehicle Perception, Intelligent Suspension Control
@@ 1 Introduction
Adjustable suspension systems, such as magnetorheological semi-active [1, 2] and fully active suspensions [3], enhance ride comfort and handling stability by enabling real-time damping control, while fully active systems can additionally adjust stiffness to adapt to varying road conditions [4]. Conventional control strategies, typically based on suspension deflection and sprung-mass acceleration feedback, improve vibration suppression to some extent. However, without road preview information, these approaches remain reactive, limiting the ability to anticipate disturbances and thereby constraining the full potential of advanced suspension control [5]. Among various road disturbances, speed bumps and potholes represent common yet critical anomalies that directly affect ride quality [6]. Early detection of such features enables predictive suspension control, allowing damping strategies to be adjusted in advance [7, 8, 9]. Existing methods for road anomaly detection rely on vehicle dynamics, vibration analysis, or multi-sensor fusion (e.g., accelerometers, gyroscopes, and GPS) [10, 11, 12]. While these approaches can effectively capture the vehicle’s response to road irregularities, they are inherently reactive, as they measure excitations only after the vehicle has encountered the disturbance. Consequently, they lack the ability to anticipate upcoming anomalies such as speed bumps and potholes, which fundamentally limits their applicability for predictive suspension control in real-time embedded systems. To overcome the limitations of reactive suspension control, Fig. 1 illustrates a vision-based preview strategy with adjustable damping for anticipating road irregularities such as speed bumps and potholes.
Figure 1: Schematic of preview control principle with adjustable suspension for speed bumps and potholes
@@
Recent advances in deep learning have enabled camera vision to emerge as an efficient and informative approach for proactive detection of road anomalies such as speed bumps and potholes. Compared to methods relying on acceleration signals, vision-based approaches provide earlier and more reliable detection, supplying critical preview information for real-time suspension control. Within this context, the YOLO (You Only Look Once) family has been influential in balancing detection accuracy with inference speed [13]. Bučko et al. [14] evaluated YOLOv3 for pothole detection under diverse lighting and weather conditions, showing that while Sparse R-CNN offered slightly higher robustness in extreme scenarios, YOLOv3 provided stable performance with lower computational demands, making it suitable for embedded deployment. YOLOv4 further introduced CSPDarknet and PANet for improved feature aggregation. Park et al. [15] demonstrated that YOLOv4-tiny could detect potholes in real time on resource-constrained devices, achieving high inference speed, while ASAD et al. [16] implemented a lightweight YOLOv4-tiny model on a Raspberry Pi, achieving approximately 90% accuracy at 31.76 FPS for edge-based pothole detection. YOLOv5 added anchor-free heads and auto-learned anchors to enhance detection flexibility [17], and YOLOv8 incorporated decoupled detection heads and improved data augmentation [18]. Khan et al. [19] further customized YOLOv8 for autonomous driving scenarios, optimizing it for road anomaly detection. POT-YOLO [20] extended YOLOv8 to detect complex pavement defects with high accuracy. Building on YOLOv11 [21], which introduces the C3k2 residual block and C2PSA attention for enhanced multiscale features and scalable deployment, we optimize its architecture and inference for real-time, accurate road anomaly detection on embedded suspension systems.
Despite advances in vision-based road anomaly detection, real-world deployment under dynamic driving conditions remains challenging. Small, distant targets and environmental factors, such as motion artifacts and illumination variability, reduce detection robustness, while limited computational resources constrain input resolution and model complexity. To address these issues, this work proposes a systematic approach that jointly optimizes network architecture, training strategy, and model quantization for efficient and accurate detection of speed bumps and potholes. The main contributions are summarized as follows:
  • 1)
@@
    Architecture and modules: A lightweight and efficient detection head (LEDH) is introduced to reduce the computational overhead of the P2 layer while preserving accuracy for small and distant targets. In combination with GhostConv modules and VoVGSCSPC blocks in the backbone and neck, the architecture improves feature extraction efficiency and robustness under diverse road and environmental conditions.
  • 2)
@@
    Training strategy: A comprehensive training pipeline is employed to enhance performance in dynamic scenarios. It incorporates NWD loss for precise localization, BCKD knowledge distillation for effective feature transfer, and data augmentation with Albumentations to simulate motion blur, illumination variations, and adverse weather.
  • 3)
@@
    Model quantization: The model is quantized to FP16 and optimized with TensorRT, achieving 139.5 FPS on NVIDIA Jetson AGX Xavier, thereby meeting real-time requirements for embedded road condition perception.
@@ 2.2 GhostConv Module
Under typical settings ($c_{1}\in[64,256]$, $k_{m}=3$, $k_{c}=5$), GhostConv reduces parameters and FLOPs to 48%–55% of SC, demonstrating its efficiency for embedded detection. Beyond computational savings, Fig. 3c shows that the augmented backbone enhances responses to distant small-scale anomalies, indicating an improved receptive field and richer feature representation.